Maxi wants to buy at least 15 kilos of bangus and pork for her birthday but she can spend no more than php 4000. A kilo of bangus cost php170 and a kilo of pork cost php 270. What are the possible number of kilos of bangus and pork can he buy?
Answer by ikleyn(53765)

.

Your inequalities are


      x +    y >= 15      (1)

   170x + 270y <= 4000    (2)

      x >= 0, y >= 0      (3)


where x = kilos of bangus and y = kilos of pork.


In visual form, this inequalities are shown in the figure below



    


                        Plot x + y = 15  (red line), 170x + 270y = 4000 (green line)



We consider the quadrant QI, since we have these inequalities (3).


The solution set is ABOVE the red line (including points in this line) and below the green line (including points in this line).


You may use any point from this set as a solution.


For example,  x= 10 (kilos of bangus)  and  y = 8 (kilos of pork).

              Check.  10*170 + 8*270 = 3860 (which is less than 4000).


And more than million other different solutions . . . 

Solved and explained.